A home run is hit in such a way that the baseball
Answer #1

let the speed be v, v*4/5*t = 120 m

by second equation of motion in vertical,

h = v*3/5*t - 4.9t^2

9 = 150*3/5 - 4.9t^2

4.9 t^2 = 81

t = sqrt(81/4.9)

= 4.066 s

v = 150/4.066 = 36.9 m/s

b) when it reaches wall. v = 36.9*4/5 i + 36.9*3/5 -4.9*81/4.9 j

= sqrt( (36.9*0.8)^2 + (36.9*0.6-81)^2)

= 65.85 m/s